# Shedding Light on Gravity: Black Hole Shadows and Lensing Signatures in Lorentz Gauge Theory

### v1

@@ Abstract
Recent advances, including gravitational wave detections and imaging of black hole shadows, have strongly validated general relativity. Nevertheless, ongoing cosmological observations suggest potential limitations of general relativity, spurring interest in modified theories of gravity. This study explores Lorentz gauge theory, an alternative gravitational framework offering promising solutions to longstanding conceptual issues in quantum gravity and cosmology. By analyzing black hole shadow structures and gravitational lensing effects—both weak and strong deflection regimes—we highlight unique observational signatures of Lorentz gauge gravity. Our findings provide valuable tools for future observational tests, potentially distinguishing these modified gravity models from general relativity and advancing our understanding of spacetime geometry and fundamental gravitational interactions.
@@ I Introduction
General relativity (GR), Einstein’s revolutionary theory, interprets gravity as a manifestation of curved spacetime, dynamically described by a metric tensor. Among its most extraordinary predictions is the existence of black holes—regions in spacetime where gravity is so intense that nothing, not even light, can escape Wald (1984); Misner et al. (1973); Virbhadra and Ellis (2000a, 2002); Claudel et al. (2001); Adler and Virbhadra (2022). The validity and precision of GR have been spectacularly confirmed through recent landmark observations, including gravitational wave detections by LIGO and Virgo collaborations et al. (2016, 2019a) and high-resolution imaging of black hole shadows by the Event Horizon Telescope (EHT) collaboration et al. (2019b, 2022).
Nevertheless, current cosmological observations revealing accelerated cosmic expansion and notable tensions in cosmological parameters have motivated significant interest in exploring modifications to the standard metric dynamics et al. (2012a). Such theoretical developments encompass alternative geometric frameworks and formulations of gravity, moving beyond the traditional metric-based description Sotiriou and Faraoni (2010). A critical conceptual challenge for classical GR is its inherent reliance on a non-degenerate metric; notably, the theory lacks a natural zero-metric ”ground state,” a deficiency linked to significant theoretical difficulties, particularly its non-renormalizability in quantum formulations Weinberg (1979).
@@
Expanding upon these conceptual advancements, the Lorentz gauge theory (LGT) emerges by gauging the Lorentz invariance of special relativistic parameterized fields, identifying a canonical clock field—the ”khronon”—as fundamental in defining spacetime geometry et al. (2017, 2006). This innovative approach inherently resolves several longstanding conceptual problems of time in quantum gravity, embedding the metric construction within the chiral Lorentz connection itself. The resulting theory elegantly integrates matter fields, naturally including spinor fields and Yang-Mills interactions, without extraneous constraints et al. (2013). Recent Hamiltonian analyses have identified that only the purely chiral curvature action extends effectively to GR, whereas the action incorporating total curvature represents a topological theory devoid of local degrees of freedom Smolin (2004). This intriguing result points towards promising pathways for quantum gravity formulations.
@@
Given these compelling theoretical motivations, further detailed studies of classical solutions are vital. Gravitational lensing phenomena, particularly weak and strong deflection angles, provide powerful observational tests for these gravity theories Weinberg (1972); Virbhadra and Ellis (2000b). The weak deflection angle regime, typically involving minor bending of light around massive objects, serves as an important testbed for GR and its alternatives Keeton and Petters (2005). Conversely, the strong deflection angle regime, encountered in proximity to compact massive objects like black holes, amplifies gravitational lensing effects, offering critical opportunities to detect subtle deviations from standard GR predictions Bozza (2002a); et al. (2012b).
@@
Additionally, the observational characterization of black hole shadows, as impressively demonstrated by the EHT’s imaging of supermassive black holes M87* and Sgr A*, has opened a novel observational window into spacetime structure and gravity theories et al. (2019b, 2022). These shadows result from photon spheres—regions where gravitational fields compel photons into unstable circular orbits, producing a distinctive dark silhouette against surrounding luminous emission Cunha and Herdeiro (2018); Cardoso and Pani (2019). Precision measurements of shadow size and shape thus serve as sensitive probes of gravity theories, potentially revealing deviations indicative of new physics Bambi and Freese (2009); Johannsen and Psaltis (2010a). Observations from the Event Horizon Telescope (EHT), notably imaging the shadow of supermassive black holes M87* and Sgr A*, have stimulated considerable research interest in testing gravity theories, investigating black holes’ scalar hair, and probing the no-hair theorem Vagnozzi and Visinelli (2019); Cardoso and Gualtieri (2016); Khodadi et al. (2020, 2021); Afrin et al. (2021); Johannsen and Psaltis (2010b). Additionally, recent analytical studies have extensively examined how environmental factors like plasma distribution influence black hole shadows and gravitational lensing, yielding critical implications for astrophysical observations Perlick et al. (2015); Bisnovatyi-Kogan and Tsupko (2010, 2018); Perlick and Tsupko (2022).
@@
Modified gravity theories, higher-dimensional models, and nonlinear electrodynamics have also been extensively explored through black hole shadow observations, constraining theoretical parameters effectively Kuang et al. (2022); Belhaj et al. (2021); Pantig et al. (2022); Puliçe et al. (2023); Allahyari et al. (2020); Okyay and Övgün (2022); Övgün (2021). Investigations into exotic objects such as naked singularities and rotating traversable wormholes further highlight the diverse phenomenological consequences visible in gravitational lensing and shadow imaging Shaikh et al. (2019); Joshi et al. (2020); Gyulchev et al. (2018).
@@ II Black holes in LGT
In this framework, the dynamical fields include a scalar field $\phi^{\alpha}$, which transforms under the fundamental representation of the Lorentz group, and a connection 1‐form $\omega^{a}{}_{b}$ in the adjoint representation. The connection naturally introduces the covariant exterior derivative $D$, and its associated curvature 2‐form is defined as $R^{a}{}_{b}=d\omega^{a}{}_{b}+\omega^{a}{}_{c}\wedge\omega^{c}{}_{b}.$ This expression encapsulates the nontrivial geometry of the SO(1,3) principal bundle over the 4-dimensional manifold $M$ Koivisto (2023); Koivisto and Zheng (2025); Wiesendanger (2019). In essence, the curvature $R^{a}{}_{b}$ measures the failure of the connection to be locally pure gauge, and it plays a central role in the gravitational dynamics of the theory.
For vacuum solutions, the field equations simplify considerably, leading to explicit forms for the metric functions. In particular, the lapse function $f(r)$, which can be interpreted as the gravitational potential, is given by Koivisto and Zheng (2025)
$f(r)=\frac{1}{A^{2}_{0}}-\frac{m_{S}}{4\pi m^{2}_{P}r}\,,$ (1)
where $m_{P}$ and $m_{S}$ are mass parameters, and $A_{0}$ is the connection coefficient. This equation not only determines the redshift of clocks (through the time component of the metric) but also implicitly affects the spatial geometry by setting a scale for the curvature. Notice that the first term, $1/A_{0}^{2}$, represents a constant offset that directly modifies the gravitational potential, while the second term shows the familiar $1/r$ falloff modulated by the mass parameters.
Substituting this lapse function into the general static, spherically symmetric metric ansatz yields
$ds^{2}=-\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)dt^{2}+\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)^{-1}dr^{2}$
$+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (2)
as demonstrated in Koivisto and Zheng (2025). This metric clearly illustrates how the parameter $A_{0}$ not only scales the time component but also rescales the radial part of the geometry. In effect, the connection coefficient $A_{0}$ plays a dual role, modulating the gravitational redshift and the spatial curvature simultaneously.
@@
A particularly illuminating case arises when one sets $A_{0}=\pm 1$. In this situation, the metric reduces to the familiar Schwarzschild solution, thereby recovering the standard event horizon at $r=2M$ (after redefining the combination $m_{S}/(4\pi m_{P}^{2})$ as $2M$, with $M$ representing the mass of the black hole). This equivalence underscores that $A_{0}$ acts as a scaling parameter: deviations from $|A_{0}|=1$ yield modified spacetimes. Specifically, for $A_{0}>1$, the effective radius of the event horizon increases, and the lapse function $f(r)$ decreases, suggesting that the black hole appears “inflated” compared to its Schwarzschild counterpart. Conversely, if $A_{0}<1$, the event horizon contracts, potentially leading to radical changes in the black hole’s structure and, in extreme cases, to the complete disappearance of the horizon.
The equations not only define the geometric structure of the theory but also highlight the significant role played by the connection parameter $A_{0}$ in determining the physical properties of static, spherically symmetric spacetimes. The interplay between the gravitational potential (via $f(r)$) and the scaling parameter $A_{0}$ provides deeper insights into how modifications in the connection can lead to deviations from classical solutions, such as the Schwarzschild metric.
The work in Abdelqader and Lake (2015) introduces a set of curvature scalars designed to identify key geometric features of black hole spacetimes—such as the event horizon, the ergosurface—and to provide measures for physical properties like mass and spin Abdelqader and Lake (2015); Tavlayan and Tekin (2020). These scalars are constructed from the Weyl tensor $C_{\mu\nu\alpha\beta}$ and its left dual ${}^{*}C_{\mu\nu\alpha\beta}$, as well as their covariant derivatives. The definitions are as follows:
@@
These invariants are powerful because, as shown in Tavlayan and Tekin (2020), one may locate the event horizon of Schwarzschild-like black holes by finding where $I_{3}$ vanishes. In these spacetimes, $I_{3}$ is positive outside the horizon, becomes zero precisely at the horizon, and is negative inside.
To illustrate these ideas in a simpler setting, consider the induced metric on a two-dimensional hypersurface with coordinates $(t,r)$. In the induced coordinates, the metric takes the diagonal form
$\gamma_{ij}=\left(\begin{array}[]{cc}f(r)&0\\ 0&\frac{1}{f(r)}.\end{array}\right)$ (5)
where $f(r)$ is the lapse function. This metric form is particularly useful when analyzing horizon properties since it retains the essential radial dependence.
An important quantity computed from the induced metric is the Kretschmann invariant, which for this two-dimensional geometry is given by Pantig and Övgün (2022)
$K={}^{\Sigma}I_{1}={}^{\Sigma}R_{ijkl}{}^{\Sigma}R^{ijkl}=\left({\frac{{\rm d}^{2}}{{\rm d}{r}^{2}}}f(r)\right)^{2},$ (6)
@@
which also vanishes at the horizon. These expressions underline how derivatives of the lapse function encode critical geometric transitions associated with the horizon.
Next, by inserting the lapse function from Eq. (1) into the induced metric Eq. (5), we can compute explicit forms for these invariants. The Kretschmann invariant associated with the induced metric becomes
@@
Thus, the invariants derived from the induced metric not only detect the presence of the event horizon but also give a clear algebraic relation for its location in terms of the parameters $A_{0}$, $m_{S}$, and $m_{P}$.
Finally, the Kretschmann scalar for the full four-dimensional spacetime is given by

### v2

@@ Abstract
Recent advances, including gravitational wave detections and imaging of black hole shadows, have strongly validated general relativity. Nevertheless, ongoing cosmological observations suggest potential limitations of general relativity, spurring interest in modified theories of gravity. This work investigates the Lorentz-gauge formulation of gravity-a novel framework that addresses key conceptual challenges in quantum gravity and cosmology by leveraging the recent black‐hole solutions presented in [10.1103/PhysRevD.111.064008] Koivisto and Zheng (2025). By analyzing black hole shadow structures and gravitational lensing effects—both weak and strong deflection regimes—we highlight unique observational signatures of Lorentz gauge gravity. Our findings provide valuable tools for future observational tests, potentially distinguishing these modified gravity models from general relativity and advancing our understanding of spacetime geometry and fundamental gravitational interactions.
@@ I Introduction
General relativity (GR), Einstein’s revolutionary theory, interprets gravity as a manifestation of curved spacetime, dynamically described by a metric tensor. Among its most extraordinary predictions is the existence of black holes—regions in spacetime where gravity is so intense that nothing, not even light, can escape Wald (1984); Misner et al. (1973); Virbhadra and Ellis (2000, 2002); Claudel et al. (2001); Adler and Virbhadra (2022). The validity and precision of GR have been spectacularly confirmed through recent landmark observations, including gravitational wave detections by LIGO and Virgo collaborations Abbott et al. (2016, 2019) and high-resolution imaging of black hole shadows by the Event Horizon Telescope (EHT) collaboration [The Event Horizon Telescope Collaboration] et al. (2019a, b, 2021, 2022).
The recent advent of very‐long‐baseline interferometry (VLBI) at millimeter and sub‐millimeter wavelengths has opened an unprecedented window into the strong‐field regime of gravity near supermassive black holes. Early proposals to image the shadow of SgrA* using space–ground VLBI were pioneered by Zakharov et. al, who demonstrated that RADIOASTRON could in principle measure black‐hole parameters from the Galactic center shadow Zakharov et al. (2005). Subsequent theoretical and numerical studies have refined this concept, showing how the shadow can be reconstructed and used as a precision test of general relativity (GR) and its alternatives Zakharov (2024). In parallel, monitoring of stellar orbits—most notably the relativistic motion of the S2 star-provides complementary tests of gravity in the same region Bambhaniya et al. (2024). On the theoretical side, analytic expressions for the shadow radius in charged spacetimes have been derived for both Reissner-Nordström and tidal‐charge metrics. Direct measurements of black‐hole charge parameters were first explored in Zakharov et al. (2005), and later generalized to include higher‐dimensional and dilatonic extensions Zakharov et al. (2012); Zakharov (2014). More recent work has exploited the EHT observations to constrain both electric and tidal charges in M87 and SgrA* shadows Zakharov (2022a, b). These studies underline the power of shadow measurements to place stringent bounds on deviations from the Kerr geometry. Despite the close relation between photon spheres and shadow boundaries in Kerr and Reissner–Nordström spacetimes, it has been shown that the existence of circular photon orbits alone does not guarantee an observable shadow in more general spherically symmetric metrics Zakharov (2025a, b). This distinction highlights the necessity of a careful analysis of the causal structure and lensing properties within any modified‐gravity framework. Finally, comprehensive reviews of the Galactic Centre as a laboratory for gravity and dark‐matter theories de Laurentis et al. (2023) and of the current status and future prospects of black‐hole imaging experiments Genzel et al. (2024) provide essential context for ongoing and forthcoming observations. Recent EHT images have been used to place stringent observational bounds on a variety of modified gravity theories, excluding compact objects in baseline mimetic gravity Khodadi et al. (2024), constraining standard‐model extension coefficients in Lorentz‐violating frameworks Khodadi and Lambiase (2022), and testing Einstein‐Æther gravity against the shadow of M87* Khodadi and Saridakis (2021). In this work, we build upon these theoretical foundations and observational breakthroughs to examine the shadows and lensing signatures arising in Lorentz Gauge Theory (LGT).
Nevertheless, current cosmological observations revealing accelerated cosmic expansion and notable tensions in cosmological parameters have motivated significant interest in exploring modifications to the standard metric dynamics Clifton et al. (2012). Such theoretical developments encompass alternative geometric frameworks and formulations of gravity, moving beyond the traditional metric-based description Sotiriou and Faraoni (2010). A critical conceptual challenge for classical GR is its inherent reliance on a non-degenerate metric; notably, the theory lacks a natural zero-metric ”ground state,” a deficiency linked to significant theoretical difficulties, particularly its non-renormalizability in quantum formulations Weinberg (1979).
@@
Expanding upon these conceptual advancements, the LGT emerges by gauging the Lorentz invariance of special relativistic parameterized fields, identifying a canonical clock field fundamental in defining spacetime geometry Złośnik and Westman (2017); Zlosnik et al. (2006); Złośnik et al. (2018); Koivisto and Zlosnik (2023); Nikjoo and Zlosnik (2024). This innovative approach inherently resolves several longstanding conceptual problems of time in quantum gravity, embedding the metric construction within the chiral Lorentz connection itself. The resulting theory elegantly integrates matter fields, naturally including spinor fields and Yang-Mills interactions, without extraneous constraints Alexandrov (2014). Recent Hamiltonian analyses have identified that only the purely chiral curvature action extends effectively to GR, whereas the action incorporating total curvature represents a topological theory devoid of local degrees of freedom Smolin (2004). This intriguing result points towards promising pathways for quantum gravity formulations.
The primary motivation for investigating LGT stems from its attempt to address several foundational problems in gravitational theory from a ”pregeometric” standpoint. Unlike GR, where the metric is a fundamental field postulated a priori, LGT begins with more basic entities: a gauge connection for the Lorentz group and a scalar field. The spacetime metric, and thus the entire geometric framework of GR, is not fundamental but rather an emergent, composite structure that arises in a symmetry-broken phase of the theory.
This approach offers several distinct advantages and potential solutions to long-standing issues: Addressing the Metric’s Origin and Non-Renormalisability: By treating spacetime geometry as emergent, LGT directly confronts the question of why a metric exists at all. The theory possesses a ”topological phase” where no local gravitational degrees of freedom propagate. This phase is described by a topological field theory which is expected to be more manageable at the quantum level. GR, with its dynamical degrees of freedom, is then proposed to emerge from this simpler, topological foundation via a mechanism of symmetry breaking Mielke (2017). This provides a conceptual pathway towards a quantum theory of gravity that sidesteps the problems associated with the non-renormalisability of the metric in standard quantisation approaches. A Dynamical Solution to the ”Problem of Time”: The scalar field $\phi^{a}$ in the fundamental representation of the Lorentz group is not an arbitrary addition. As originally proposed in Złośnik et al. (2018), this field can be identified with a canonical ”clock,” which we term the khronon in its time-like phase. Its gradient, $\bm{\mathrm{e}}^{a}=\bm{\mathrm{D}}\phi^{a}$, defines the emergent coframe. This mechanism dynamically introduces a physical reference frame into the theory. Consequently, it provides a natural, built-in resolution to the ”problem of time” that plagues canonical quantum gravity, without needing to postulate external reference matter fields like the Brown-Kuchar dust Brown and Kuchar (1995). In our framework, such a dust-like energy component emerges naturally from the theory’s structure. Natural Coupling to Matter: The first-order, gauge-theoretic nature of LGT allows for a remarkably natural and consistent coupling to matter fields. Spinor fields and other Yang-Mills gauge fields can be incorporated into the pregeometric framework in a way that is often more straightforward than in other formulations of GR or its modifications Gallagher et al. (2024). Connection to Yang-Mills and Topological Field Theories: The action for LGT can be formulated as a ”curvature-squared” theory, drawing a direct and powerful analogy with the successful Yang-Mills theories of the Standard Model. This structural similarity, combined with its origin as a chiral projection of a topological field theory Nikjoo and Zlosnik (2024), places LGT on a solid theoretical foundation and opens promising avenues for its quantization.
Therefore, Lorentz Gauge Gravity is not merely another alternative to GR, but a comprehensive framework that aims to reconstruct gravity from more fundamental gauge principles. It offers potential resolutions to deep conceptual problems while maintaining a close structural relationship with other well-understood field theories. This motivates our detailed investigation into its classical solutions, such as the black hole spacetimes explored in this work.
@@
Given these compelling theoretical motivations, further detailed studies of classical solutions are vital. Gravitational lensing phenomena, particularly weak and strong deflection angles, provide powerful observational tests for these gravity theories Weinberg (1972); Virbhadra and Ellis (2000). The weak deflection angle regime, typically involving minor bending of light around massive objects, serves as an important testbed for GR and its alternatives Keeton and Petters (2005). Conversely, the strong deflection angle regime, encountered in proximity to compact massive objects like black holes, amplifies gravitational lensing effects, offering critical opportunities to detect subtle deviations from standard GR predictions Bozza (2002); Tsukamoto (2018a). Gravitational lensing effects around black holes have been extensively studied, incorporating corrections from topological charges, holonomy, loop quantum gravity, and polymerized models, as explored in various theoretical frameworks Soares et al. (2024, 2023a, 2023b, 2025); Kumar Walia (2023).
@@
Additionally, the observational characterization of black hole shadows, as impressively demonstrated by the EHT’s imaging of supermassive black holes M87* and Sgr A* [The Event Horizon Telescope Collaboration] et al. (2019a, 2022), has opened a novel observational window into spacetime structure and gravity theories. These shadows result from photon spheres—regions where gravitational fields compel photons into unstable circular orbits, producing a distinctive dark silhouette against surrounding luminous emission Cunha and Herdeiro (2018); Cardoso and Pani (2019). Precision measurements of shadow size and shape thus serve as sensitive probes of gravity theories, potentially revealing deviations indicative of new physics Bambi and Freese (2009); Johannsen and Psaltis (2010). Observations from the EHT, notably imaging the shadow of supermassive black holes M87* and Sgr A*, have stimulated considerable research interest in testing gravity theories, investigating black holes’ scalar hair, and probing the no-hair theorem Vagnozzi and Visinelli (2019); Cardoso and Gualtieri (2016); Khodadi et al. (2020, 2021); Afrin et al. (2021); Johannsen and Psaltis (2010). Additionally, recent analytical studies have extensively examined how environmental factors like plasma distribution influence black hole shadows and gravitational lensing, yielding critical implications for astrophysical observations Perlick et al. (2015); Bisnovatyi-Kogan and Tsupko (2010, 2018); Perlick and Tsupko (2022). Recent studies have explored the optical properties and shadows of various black hole models, revealing distinct holographic and shadow images influenced by factors such as Lorentz symmetry breaking, rotation, and surrounding matter Zeng et al. (2025); Yang et al. (2025); He et al. (2024, 2025); Li et al. (2020); Hu et al. (2021); Guo and Li (2020).
@@
Modified gravity theories, higher-dimensional models, and nonlinear electrodynamics have also been extensively explored through black hole shadow observations, constraining theoretical parameters effectively Allahyari et al. (2020); Belhaj et al. (2021); Khodadi and Saridakis (2021); Okyay and Övgün (2022); Övgün (2021); Kuang et al. (2022); Khodadi and Lambiase (2022); Pantig et al. (2022); Puliçe et al. (2023); Khodadi et al. (2024). Investigations into exotic objects such as naked singularities and rotating traversable wormholes further highlight the diverse phenomenological consequences visible in gravitational lensing and shadow imaging Shaikh et al. (2019); Joshi et al. (2020); Gyulchev et al. (2018).
@@ II Black holes in LGT
This section outlines a theoretical model based on Lorentz gauge theory Koivisto and Zheng (2025) . The fundamental components of this theory are a dimensionless scalar field, denoted as $\phi^{a}$, which exists in the fundamental representation of the Lorentz group, and the canonical connection 1-form, $\bm{\omega}^{a}{}_{b}$, in the adjoint representation. This connection facilitates the definition of a covariant exterior derivative, $\bm{\mathrm{D}}$. The curvature of this connection is expressed as Koivisto and Zheng (2025) $\bm{R}^{a}{}_{b}=\bm{\mathrm{d}}\bm{\omega}^{a}{}_{b}+\bm{\omega}^{a}{}_{c}\wedge\bm{\omega}^{c}{}_{b}$. The entire framework is established on an SO(1,3) bundle over a four-dimensional manifold, $M$.
The action of the theory is constructed as an integral of a polynomial 4-form built from the available fields. When considering only the connection, the most general action is a topological field theory with two parameters, $g_{1}$ and $g_{2}$:
$I_{(0)}=\int_{M}\left(g_{1}\eta_{ab}\eta^{cd}+g_{2}\epsilon_{ab}{}^{cd}\right)\bm{R}^{a}{}_{c}\wedge\bm{R}^{b}{}_{d}$ (1)
Here, $\eta_{ab}$ and $\epsilon_{ab}{}^{cd}$ represent the invariants of the Lorentz algebra. This action, $I_{(0)}$, describes a topological field theory, which lacks local degrees of freedom in the bulk manifold. Such theories are mathematically significant and offer a potential pathway to quantum gravity, where a theory like General Relativity (GR) could emerge through a process of symmetry breaking.
To introduce dynamics, the scalar field $\phi^{a}$ is incorporated. The construction of the action is guided by a crucial principle: a global shift symmetry. This symmetry dictates that the theory should only depend on the covariant changes in $\phi^{a}$, rendering its absolute value physically meaningless. A transformation of the form $\phi^{a}\rightarrow\phi^{a}+\xi^{a}$ (where $\bm{\mathrm{D}}\xi^{a}=0$) must leave the action unchanged. Under this constraint, three additional independent 4-forms can be added to the action:
$I_{(4)}=\lambda\int_{M}\epsilon_{abcd}\bm{\mathrm{D}}\phi^{a}\wedge\bm{\mathrm{D}}\phi^{b}\wedge\bm{\mathrm{D}}\phi^{c}\wedge\bm{\mathrm{D}}\phi^{d}$ (2)
and
$I_{(2)}=g_{3}\int_{M}\bm{\mathrm{D}}\phi^{a}\wedge\bm{\mathrm{D}}\phi^{b}\wedge\bm{R}_{ab}+g_{4}\int_{M}\epsilon_{abcd}\bm{\mathrm{D}}\phi^{a}\wedge\bm{\mathrm{D}}\phi^{b}\wedge\bm{R}^{cd}$ (3)
If the shift symmetry is not held as exact, and the action is permitted to change by a boundary term, additional terms ($\tilde{I}_{(2)}$ and $\tilde{I}_{(4)}$) can be considered. These terms are equivalent to $I_{(2)}$ and $I_{(4)}$ up to boundary terms and suggest that the theory’s dynamics can be viewed as emerging from the topological theory through a deformation of the inner product.
Further analysis reveals that the combined action $I=I_{(0)}+I_{(2)}+I_{(4)}$ remains devoid of local dynamics under the specific condition that the coupling constant $g_{3}=0$. This is due to an enhanced symmetry in the theory.
To obtain a desirable extension of General Relativity, this symmetry must be broken. This is achieved by selecting specific values for the couplings $g_{3}$ and $g_{4}$ that correspond to a maximally chiral projection of the curvature, denoted as $\bm{R}^{+}_{ab}$. The resulting action is:
$I_{+}=\int_{M}\bm{\mathrm{D}}\phi^{a}\wedge\bm{\mathrm{D}}\phi^{b}\wedge\bm{R}^{+}_{ab}$ (4)
The explicit form of this chiral projection, $\bm{R}^{+a}{}_{b}=\frac{1}{2}\left(\delta^{a}_{c}\delta^{d}_{b}-\frac{i}{2}\epsilon^{ad}{}_{bc}\right)\bm{R}^{c}{}_{d}$, necessitates a complexification of the Lorentz group. While this review adheres to the conventional metric signature, it is noted that a Euclidean formulation might offer a more fundamental perspective, particularly concerning the symmetries of the Standard Model and insights from twistor geometry.
In conclusion, the full theory, described by the action $I=I_{(0)}+I_{(2)}+I_{(4)}$, is governed by five dimensionless coupling constants. The theory transitions from being topologically trivial to dynamically active by adjusting the parameter $g_{3}$. By setting $g_{4}=-i/4$ and treating $g_{3}=a/2$ as an interpolating parameter, the quadratic part of the action becomes:
$I_{(2)}=\frac{1}{2}\int_{M}\bm{\mathrm{D}}\phi_{a}\wedge\bm{\mathrm{D}}\phi^{b}\wedge\left(a\delta^{a}_{c}\delta^{d}_{b}-\frac{i}{2}\epsilon^{ad}{}_{bc}\right)\bm{R}^{c}{}_{d}$ (5)
Here, the parameter $a$ (termed the ”dymaxion”) can be conceptualized as a field that drives the transition between phases: $a=0$ corresponds to the topological phase, while $a=1$ represents the fully dynamical, extended chiral GR phase. This framework provides a mechanism for the emergence of a dynamical spacetime from an underlying topological structure. For the remainder of this work, the theory is analyzed in its dynamical phase where $a=1$.
@@
In this framework, the dynamical fields include a scalar field $\phi^{\alpha}$, which transforms under the fundamental representation of the Lorentz group, and a connection 1‐form $\omega^{a}{}_{b}$ in the adjoint representation. The connection naturally introduces the covariant exterior derivative $D$, and its associated curvature 2‐form is defined as $R^{a}{}_{b}=d\omega^{a}{}_{b}+\omega^{a}{}_{c}\wedge\omega^{c}{}_{b}$. This expression encapsulates the nontrivial geometry of the SO(1,3) principal bundle over the 4-dimensional manifold $M$ Koivisto (2023); Koivisto and Zheng (2025); Wiesendanger (2019). In essence, the curvature $R^{a}{}_{b}$ measures the failure of the connection to be locally pure gauge, and it plays a central role in the gravitational dynamics of the theory. Within LGT, static and spherically symmetric spacetimes are characterized by metric functions $f(r)$ and $g(r)$, which are determined by solving the field equations. An interesting feature of these equations is the appearance of an extra parameter $A_{0}$, which originates from the structure of the connection and reflects the choice of the observer’s frame of reference. This parameter is crucial since it adjusts the scale of both the temporal and spatial parts of the metric. For vacuum solutions, the field equations simplify considerably, leading to explicit forms for the metric functions Koivisto and Zheng (2025)
$ds^{2}=-\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)dt^{2}+\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (6)
@@
where $m_{P}$ and $m_{S}$ are mass parameters, and $A_{0}$ is the connection coefficient. This metric clearly illustrates how the parameter $A_{0}$ not only scales the time component but also rescales the radial part of the geometry. In effect, the connection coefficient $A_{0}$ plays a dual role, modulating the gravitational redshift and the spatial curvature simultaneously. A particularly illuminating case arises when one sets $A_{0}=\pm 1$. In this situation, the metric reduces to the familiar Schwarzschild solution, thereby recovering the standard event horizon at $r=2M$ (after redefining the combination $m_{S}/(4\pi m_{P}^{2})$ as $2M$, with $M$ representing the mass of the black hole). This equivalence underscores that $A_{0}$ acts as a scaling parameter: deviations from $|A_{0}|=1$ yield modified spacetimes. Specifically, for $A_{0}>1$, the effective radius of the event horizon increases, and the lapse function $f(r)$ decreases, suggesting that the black hole appears “inflated” compared to its Schwarzschild counterpart. Conversely, if $A_{0}<1$, the event horizon contracts, potentially leading to radical changes in the black hole’s structure and, in extreme cases, to the complete disappearance of the horizon. The equations not only define the geometric structure of the theory but also highlight the significant role played by the connection parameter $A_{0}$ in determining the physical properties of static, spherically symmetric spacetimes. The interplay between the gravitational potential (via $f(r)$) and the scaling parameter $A_{0}$ provides deeper insights into how modifications in the connection can lead to deviations from classical solutions, such as the Schwarzschild metric. The work in Abdelqader and Lake (2015) introduces a set of curvature scalars designed to identify key geometric features of black hole spacetimes—such as the event horizon, the ergosurface—and to provide measures for physical properties like mass and spin Abdelqader and Lake (2015); Tavlayan and Tekin (2020). These scalars are constructed from the Weyl tensor $C_{\mu\nu\alpha\beta}$ and its left dual $^{*}C_{\mu\nu\alpha\beta}$, as well as their covariant derivatives. The definitions are as follows:
@@
These invariants are powerful because, as shown in Tavlayan and Tekin (2020), one may locate the event horizon of Schwarzschild-like black holes by finding where $I_{3}$ vanishes. In these spacetimes, $I_{3}$ is positive outside the horizon, becomes zero precisely at the horizon, and is negative inside. To illustrate these ideas in a simpler setting, consider the induced metric on a two-dimensional hypersurface with coordinates $(t,r)$. In the induced coordinates, the metric takes the diagonal form
$\gamma_{ij}=\left(\begin{array}[]{cc}f(r)&0\\ 0&\frac{1}{f(r)}\end{array}\right).$ (9)
where $f(r)$ is the lapse function. This metric form is particularly useful when analyzing horizon properties since it retains the essential radial dependence. An important quantity computed from the induced metric is the Kretschmann invariant, which for this two-dimensional geometry is given by Pantig and Övgün (2022)
$K={}^{\Sigma}I_{1}={}^{\Sigma}R_{ijkl}{}^{\Sigma}R^{ijkl}=\left({\frac{{\rm d}^{2}}{{\rm d}{r}^{2}}}f(r)\right)^{2}.$ (10)
@@
which also vanishes at the horizon. These expressions underline how derivatives of the lapse function encode critical geometric transitions associated with the horizon. Next, by inserting the lapse function from Eq. (6) into the induced metric Eq. (9), we can compute explicit forms for these invariants. The Kretschmann invariant associated with the induced metric becomes
@@
Thus, the invariants derived from the induced metric not only detect the presence of the event horizon but also give a clear algebraic relation for its location in terms of the parameters $A_{0}$, $m_{S}$, and $m_{P}$. Finally, the Kretschmann scalar for the full four-dimensional spacetime is given by